OpenStudy (anonymous):

What is the approximate perimeter of a semicircle with a radius of 6 cm? Use pi star times pi= 3.14. Express your final answer to the nearest hundredth.

OpenStudy (anonymous):

By any chance, what's the perimeter (circumference) of a circle? Do you know? ^^

OpenStudy (anonymous):

ya

OpenStudy (igreen):

Okay, what is it?

OpenStudy (anonymous):

c = pi*d

OpenStudy (igreen):

Yes..or C = 2 * pi * r

OpenStudy (anonymous):

Right... but what if it's radius you're given?

OpenStudy (anonymous):

then u just multiply it by 2 in ur head .-.

OpenStudy (igreen):

Lol

OpenStudy (igreen):

Just plug in what you know..the radius and 3.14 for pi..

OpenStudy (igreen):

Find it and then divide by 2, because it's a semi-circle.

OpenStudy (anonymous):

Of course. Now,we have a semicircle with radius 6. If it was a circle instead, what would its perimeter (circumference) be?

OpenStudy (anonymous):

right but its semi

OpenStudy (igreen):

Yes, so just find it and divide by 2..

OpenStudy (igreen):

\(C = 2 \pi r\) \(C =~?\)

OpenStudy (igreen):

More like \(C = \dfrac{2 \pi r}{2}\)

OpenStudy (anonymous):

18.84

OpenStudy (igreen):

Correct, that's our answer.